You found a treasure map and it said: start at the well, go 100 m straight South, then 30 m West, then 25 straight East, and finally 45 m, North. Using the component method of vector addition, how far from the well and in what direction is the hidden treasure?


1
Expert's answer
2021-02-19T10:28:56-0500

"\text{let }\vec{X}(x,y) \text{ and }(x,y)\text{ vector component:}"

"\text{where }x=x*\vec{n};y=x*\vec{e}"

"\vec{n}-\text{unit vector pointing North}"

"\vec{e}-\text{unit vector pointing East}"

"\text{negative x component values South direction}"

"\text{negative н component values West direction}"

"\text{go 100 m straight South}-(-100,0)"

"\text{then 30 m West}-(0,-30)"

"\text{then 25 straight East}-(0,25)"

"\text{and finally 45 m, North} -(45,0)"

"\text{point of treasure}-(-100+0+45,0-30+25+0);"

"(-55,-5)"

"S= \sqrt{(-55)^2+(-5)^2}\approx52.23"

"cotg(\alpha)=\frac{55}{5}=11;\alpha\approx5.2\degree"

"\text{Answer: }\text{treasure is located 52.23 m }5.2\degree\text{southwestward}"